# Random Utility with Aggregation

### v2

@@ 2.2 Datasets and RU-rationalities
Let $\mathscr{D}\subseteq 2^{\mathscr{A}}\setminus\{\emptyset\}$ be the set of observable choice sets. The dataset in this paper is a stochastic choice function $\rho$ on $\mathscr{D}$. That is, the dataset is a function $\rho:\mathscr{D}\times\mathscr{A}\rightarrow\mathbb{R}_{+}$ such that for any $A\in\mathscr{D}$, $\sum_{a\in A}\rho(A,a)=1$, $\rho(A,a)\geq 0$ for all $a\in A$ and $\rho(A,a)=0$ for all $a\notin A$. Note that the stochastic choice function is defined over menus of categories in $\mathscr{A}$}, not over the underlying alternatives $\mathscr{X}$. Thus, we sometimes call $\rho$ reduced dataset.
@@
For simplicity, we assume that $\rho$ is defined on all subsets of $\mathscr{A}$ (i.e., ${\cal D}=2^{\mathscr{A}}\setminus\{\emptyset\}$). However this assumption is not necessary for our main results. After we present our theorems, we provide a remark in which we specify a weaker domain assumption. (See Remark 3.4 for Theorem 3.1, Remark 3.11 for Theorem 3.6, and Remark 3.13 for Theorem 3.12.)
As mentioned in the introduction, in discrete choice analysis, the standard method in the presence of aggregated categories is to assume aggregated random utility models (ARUMs), which is a probability distribution over rankings over $\mathscr{A}$.
@@
In our framework, preferences are defined on the set $\mathscr{X}$ of underlying primitive alternatives rather than on their aggregates. The components of an aggregate category may differ substantially—for example, across versions of the RAV4 or within a heterogeneous outside option—and these differences matter to agents. Accordingly, the relevant random utility model should be formulated over $\mathscr{X}$.
To connect these two levels of analysis, we introduce a formal mapping called a composition distribution which links the aggregated categories to the underlying alternatives in their composition. Given an aggregation function, this mapping $\lambda$ captures the distribution of available underlying alternatives faced by agents in the population, which may not be observable to the analyst.
@@
Given an aggregation function $X$, a set of functions $(\lambda_{A})_{A\subseteq\mathscr{A}}$ is called a composition distribution if for any $A\in\mathscr{D}$,
@@
The value of the composition distribution $\lambda_{A}((S_{a})_{a\in A})$ represents the frequency with which each aggregated category $a$ is composed of the subset $S_{a}\subseteq X(a)$ within the choice set $A\subseteq\mathscr{A}$. We allow $\lambda$ to depend on the choice set $A$, since the availability of underlying alternatives may vary across markets—for example, it may depend on which stores consumers can access. In Example 2.1, eggs may be unavailable in one market (e.g., due to a bird flu outbreak), making omelettes infeasible, while pancake ingredients remain widely available. In this case, $\lambda_{A}$ would assign higher probability to $S_{\text{out}}$ containing pancakes than to sets including omelettes.
We now define RU-rationality in the presence of aggregated categories whose compositions are unknown to the analyst. To motivate the definition, return to the cereal example and consider a distribution $\mu_{{\scriptscriptstyle\mathscr{X}}}$ over the underlying alternatives (the cereals and other breakfast options). Let $D$ be the set of cereals available in a store and let $x\in D$ denote one such cereal. For a subpopulation of consumers whose outside option consists only of omelettes and pancakes, the market share of $x$ is $\mu_{{\scriptscriptstyle\mathscr{X}}}(\succ\mid x\succ y\text{ for all }y\in D\cup\{\text{omelette},\text{pancakes}\})$. Analogous expressions arise for other possible compositions of the outside option. Aggregating across these compositions using $\lambda_{A}$ yields equation (1) in the introduction. The following extends this construction to the case with multiple aggregated categories:
@@
A stochastic choice function $\rho$ is RU-rational if there exist $\mu_{{\scriptscriptstyle\mathscr{X}}}\in\Delta(\mathscr{L}(\mathscr{X}))$, an aggregation function $X$, and a composition distribution $\lambda$ such that for any $a\in A\in\mathscr{D}$,
$\rho(A,a)=\sum_{(S_{b})_{b\in A}\in\underset{b\in A}{\prod}(2^{X(b)}\setminus\{\emptyset\})}\lambda_{A}((S_{b})_{b\in A})\mu_{{\scriptscriptstyle\mathscr{X}}}(\succ\in\mathscr{L}(\mathscr{X})|\exists x\in S_{a},\ \forall y\in\underset{b\in A\setminus a}{\cup}S_{b},\ x\succ y).$ (3)
We say that the pair $(\mu_{{\scriptscriptstyle\mathscr{X}}},\lambda)$ rationalizes $\rho$. [^9] Strictly speaking, we should write $(\mu_{{\scriptscriptstyle\mathscr{X}}},X,\lambda)$ rationalizes $\rho$. However, since the role of $X$ is implicit in $\lambda$, we will simply write $(\mu_{{\scriptscriptstyle\mathscr{X}}},\lambda)$.
Note that the term $\mu_{{\scriptscriptstyle\mathscr{X}}}(\cdots)$ on the right-hand side of equation (3) represents the total market share of category $a$ from menu $A$ within the subpopulation of consumers for whom each aggregated alternative $b$ corresponds to the subset $S_{b}\subseteq X(b)$. The term $\lambda_{A}((S_{b})_{b\in A})$ denotes the proportion of such consumers in the overall population. Thus, the right-hand side of equation (3) is the weighted average of these market shares, with weights given by $\lambda$.
@@
Definition 2.5 assumes independence between the distribution $\mu_{{\scriptscriptstyle\mathscr{X}}}$ over preference rankings and the composition distribution $\lambda$ over underlying alternatives. While one could consider a more general framework allowing dependence between these components, such a model may lack testable implications, since the joint distribution could vary with the menu as $\lambda$ does. In particular, if the agent’s preferences may depend on composition, such dependence would allow the distribution over rankings to vary with the menu, rendering the model unfalsifiable. If, however, we only allow the composition distribution $\lambda$ to depend on the preferences, the characterization remains the same. See Remark 3.3 for more details.
@@
ARU-rationality implies RU-rationality. To see why suppose $\rho$ is rationalized by $\mu_{{\scriptscriptstyle\mathscr{A}}}$ over categories $\mathscr{A}$. Choose an aggregation function $X$ so that $X(a)$ is a singleton for all $a\in\mathscr{A}$. Label the sole element of $X(a)$ by $x_{a}$. Then for each $\succ$ in the support of $\mu_{{\scriptscriptstyle\mathscr{A}}}$ define $\succ^{\prime}$ by $x_{a}\succ^{\prime}x_{b}$ whenever $a\succ b$ and define all other comparisons arbitrary. By setting $\mu_{{\scriptscriptstyle\mathscr{X}}}(\succ^{\prime})=\mu_{{\scriptscriptstyle\mathscr{A}}}(\succ)$ for all $\succ$ in the support of $\mu_{{\scriptscriptstyle\mathscr{A}}}$, it is apparent that $(\mu_{{\scriptscriptstyle\mathscr{X}}},\lambda)$ rationalizes $\rho$.
@@
A stochastic choice function $\rho$ is partially RU-rational if there exists $\mu\in\Delta(\mathscr{L}(\mathscr{A}_{N}))$ such that for any $a\in A\subseteq\mathscr{A}_{N}$ such that $A\in\mathscr{D}$,
$\rho(A,a)=\mu(\succ\in\mathscr{L}(\mathscr{A}_{N})|a\succ b\text{ for all }b\in A\setminus a).$ (4)
@@
RU-rationality implies partial RU-rationality. When the menu contains only non-aggregated categories, each category maps to a single primitive alternative, so the RU-representation collapses to a standard RUM on the set of non-aggregated categories, which corresponds to partial RU rationality; $\lambda$ terms disappear because there is no compositional uncertainty on such menus.
@@
As mentioned earlier, ARU-rationality is the standard assumption in much of the empirical literature. RU-rationality notion introduced in this paper is a natural refinement that accounts for the unknown composition of aggregated categories.
In the next section, we characterize RU-rationality and show precisely how it is weaker than ARU-rationality. The characterization is derived by bridging the gap between RU-rationality and partial RU-rationality.
@@ 3 Characterization
In this section, we consider the outside option setup in which there exists only one aggregated category $a_{0}\in\mathscr{A}$. Formally, we set $\mathscr{A}_{A}=\{a_{0}\}$ and $\mathscr{A}_{N}=\mathscr{A}\setminus\{a_{0}\}$. We call $a_{0}$ the outside option.
This case is useful for two reasons. First, it reflects the common modeling practice of including an outside option—an aggregated category that captures unobserved or unlisted choices—frequently used in empirical applications. Although empirical models often involve more than one aggregated category, the outside option is particularly prominent. Second, this simplified setting serves as a foundation for analyzing the more general case with multiple aggregated categories in Section 3.2.
@@
Consider the outside option setup (i.e., $\mathscr{A}_{A}=\{a_{0}\}$). A stochastic choice function $\rho$ is RU-rational if and only if $\rho$ satisfies the following two conditions:
@@
    (Limited Monotonicity): if $b\in D\subseteq\mathscr{A}_{N}$ then $\rho(D,b)\geq\rho(D\cup\{a_{0}\},b)$,
@@
The necessity of (i) and (ii) is immediate. The sufficiency part of the proof is constructive and nontrivial: given a stochatsic chioce function $\rho$ that satisfies the two conditions (i) and (ii), we need to construct an aggregation function, a composition distribution and a distribution over linear orders on $\mathscr{X}$ that satisfies equation (3). The proof is in the appendix.
The theorem makes precise the weakness of RU-rationality relative to ARU-rationality: RU-rationality is substantially weaker. To see this, note that Limited Monotonicity requires only that adding the outside option does not increase the choice frequencies of non-aggregated categories. In particular, it does not require the monotonicity condition with respect to menus containing aggregated categories. To see this, consider a choice set $D$ such that $a_{0}\in D$ and $b\notin D$. The theorem shows that under RU-rationality, it is possible to observe
@@
which violates the standard monotonicity condition. This may arise because, in the expanded set $D\cup{b}$, the outside option $a_{0}$ may become more desirable than it was in $D$. For example, in the context of cereal choice, the presence of a fancy, expensive cereal brand may signal that the market corresponds to a high-income area leading to a change in the interpretation or attractiveness of the outside good: the outside goods may contain more appealing options such as smoked salmon or quiche. By contrast, the meanings of non-aggregated categories remain the same when $a_{0}$ is added to a choice set with only non-aggregated categories; thus the limited monotonicity condition holds.
ARU-rationality, by contrast, imposes much stronger restrictions including full monotonicity and beyond. When $\rho$ is observed on all subsets of $\mathscr{A}$ (i.e., $\mathscr{D}=2^{\mathscr{A}}\setminus\{\emptyset\}$), ARU-rationality is equivalent to the non-negativity of all Block-Marschak (BM) polynomials. Among these, monotonicity is one of the simplest and most intuitive implications. However, as Theorem 3.1 shows, RU-rationality only implies monotonicity on limited choice sets.
This result highlights the substantial gap between RU-rationality and ARU-rationality. In Section 4, we examine the specific conditions on the underlying distributions $\mu$ and $\lambda$ under which RU-rational stochastic choice functions are also compatible with an ARUM.
Two more technical remarks are in order:
@@
In our discussion, we focus on Limited Monotonicity rather than on partial RU-rationality. The reason is straightforward: partial RU-rationality is simply RU-rationality restricted to menus consisting only of non-aggregated categories, and under the full domain assumption (i.e., $\mathscr{D}=2^{\mathscr{A}}\setminus\{\emptyset\}$) it can be characterized by the non-negativity of Block–Marschak (BM) polynomials defined over sets composed solely of non-aggregated categories.
@@
In our notion of RU rationality, we assume that the composition of categories and the agent’s preferences are independent. A natural extension is to allow the composition distribution to depend on the agent’s preferences. That is, a composition distribution $\lambda_{\succ,A}$ is a distribution over  $\prod_{a\in A}(2^{X(a)}\setminus\{\emptyset\})$ that may jointly depend on the available categories $A$ and the agent’s preference $\succ\in\mathscr{L}(\mathscr{X})$. For example, we may allow the case where consumers with a preference for high quality goods tend to have a higher quality outside option. [^10] We do not, however, allow the composition of the outside option to enter into the consumer’s underlying choices. This type of dependence would lead to RUM violations in the underlying choice which would lead to a model with no observable implications.
We say that $\rho$ is General RU rational if there exists $(\mu_{{\scriptscriptstyle\mathscr{X}}},\lambda_{\succ,A})$ such that,
$\rho(A,a)=\sum_{(S_{b})_{b\in A}\in\underset{b\in A}{\prod}(2^{X(b)}\setminus\{\emptyset\})}\sum_{\succ\in\mathscr{L}(\mathscr{X})}1(\exists x\in S_{a},\ \forall y\in\underset{b\in A\setminus\{a\}}{\cup}S_{b},\ x\succ y))\lambda_{\succ,A}((S_{b})_{b\in A})\mu_{{\scriptscriptstyle\mathscr{X}}}(\succ).$ (5)
@@
For the theorem, in order to make sense of Limited Monotonicity, we need the following domain assumption on which choice probabilities are observable: if $\rho$ is defined on $(D\cup\{a_{0}\},\cdot)$, then $\rho$ is defined on $(D,\cdot)$ (i.e., if $D\cup\{a_{0}\}\in\mathscr{D}$, then $D\in\mathscr{D}$). In this case, Limited Monotonicity should be defined as: for all $D\subseteq\mathscr{A}_{N}$ such that $D\cup\{a_{0}\}\in\mathscr{D}$, $\rho$ satisfies $\rho(D,a)\geq\rho(D\cup\{a_{0}\},a)$ for all $a\in D$. In the next section (Section 3.1), we omit the assumption and provide an alternative characterization of RU rationality.

### v3

@@ 2.1 Datasets and RU-rationalities
Let $\mathscr{D}\subseteq 2^{\mathscr{A}}\setminus\{\emptyset\}$ be the set of choice sets. The dataset in this paper is a stochastic choice function $\rho$ on $\mathscr{D}$. That is, the dataset is a function $\rho:\mathscr{D}\times\mathscr{A}\rightarrow\mathbf{R}_{+}$ such that for any $A\in\mathscr{D}$, $\sum_{a\in A}\rho(A,a)=1$, $\rho(A,a)\geq 0$ for all $a\in A$ and $\rho(A,a)=0$ for all $a\notin A$. Note that the stochastic choice function is defined over menus of aggregates in $\mathscr{A}$}, not over the underlying alternatives $\mathscr{X}$. Thus, we sometimes call $\rho$ reduced dataset. We define a distance function between two stochastic choice functions $\rho$ and $\rho^{\prime}$ by $\|\rho-\rho^{\prime}\|=\sqrt{\sum_{x\in D\in\mathscr{D}}[\rho(D,x)-\rho^{\prime}(D,x)]^{2}}$ by using the standard Euclidean norm.
@@
In most parts of this paper, we assume that $\rho$ is defined on all subsets of $\mathscr{A}$ (i.e., $\mathscr{D}=2^{\mathscr{A}}\setminus\{\emptyset\}$) . This domain assumptions are technical, as they only serve to facilitate our expositions. For each main reuslt, the assumption can be weakened. For details, see Remark D.1 for Theorem 3.1; Remark D.6 for Theorem 3.2; and Remark D.4 for Theorem 3.3 in the online appendix.
As mentioned in the introduction, in discrete choice analysis, the standard method in the presence of aggregation is to assume aggregated random utility models (ARUMs), which is a probability distribution over rankings over $\mathscr{A}$.
@@
As discussed in the introduction, consumers evaluate and choose among the actual goods available to them, not among analyst-defined aggregates. Thus, RUM over the underlying alternatives $\mathscr{X}$ is the natural assumption on the data generating process. The difficulty is that the components of a non-atomic aggregate may differ substantially—for example, across beef products or within a heterogeneous outside option—and that the exact composition may vary across consumers and be unobservable to the analyst. To capture this, we introduce a composition distribution which links the aggregates to the distribution over underlying alternatives in their composition. Given an aggregation correspondence, this mapping $\lambda$ specifies the distribution of available underlying alternatives faced by agents in the population.
@@
Given an aggregation correspondence $X$, a set of functions $(\lambda_{A})_{A\in\mathscr{D}}$ is called a composition distribution if for any $A\in\mathscr{D}$,
@@
When there is no danger of confusion, we simply write $\lambda$ rather than $(\lambda_{A})_{A\in\mathscr{D}}$.
The value of the composition distribution $\lambda_{A}((S_{a})_{a\in A})$ represents the frequency with which each aggregate $a$ is composed of the subset $S_{a}\subseteq X(a)$ within the choice set $A\subseteq\mathscr{A}$. We allow $\lambda$ to depend on the choice set $A$, since the availability of underlying alternatives may vary across markets—for example, it may depend on which stores consumers can access. In Example 2.1, eggs may be unavailable in one market (e.g., due to a bird flu outbreak), making omelettes infeasible, while pancake ingredients remain widely available. In this case, $\lambda_{A}$ would assign higher probability to $S_{\text{out}}$ containing pancakes than to sets including omelettes.
We now define RU-rationality in the presence of aggregates whose compositions are unknown to the analyst. To motivate the definition, return to the cereal example and consider a distribution $\mu_{{\scriptscriptstyle\mathscr{X}}}$ over preferences among the underlying alternatives (the cereals and other breakfast options). Let $D$ be the set of cereals available in a store and let $x\in D$ denote one such cereal. For a subpopulation of consumers whose outside option consists only of omelettes and pancakes, the market share of $x$ is $\mu_{{\scriptscriptstyle\mathscr{X}}}(\succ\mid x\succ y\text{ for all }y\in D\cup\{\text{omelette},\text{pancakes}\})$. Analogous expressions arise for other possible compositions of the outside option. Aggregating across these compositions using $\lambda_{A}$ yields equation (1) in the introduction. The following extends this construction to the case with multiple non-atomic aggregates:
@@
A stochastic choice function $\rho$ is RU-rational if there exist $\mu_{{\scriptscriptstyle\mathscr{X}}}\in\Delta(\mathscr{L}(\mathscr{X}))$, an aggregation correspondence $X$, and a composition distribution $\lambda$ such that for any $a\in A\in\mathscr{D}$,
$\rho(A,a)=\sum_{(S_{b})_{b\in A}\in\underset{b\in A}{\prod}(2^{X(b)}\setminus\{\emptyset\})}\lambda_{A}\big((S_{b})_{b\in A}\big)\mu_{{\scriptscriptstyle\mathscr{X}}}\big(\succ\in\mathscr{L}(\mathscr{X})|\exists x\in S_{a},\ \forall y\in\underset{b\in A\setminus a}{\cup}S_{b},\ x\succ y\big).$ (3)
In this case, we say that the pair $(\mu_{{\scriptscriptstyle\mathscr{X}}},\lambda)$ rationalizes $\rho$. [^9] Strictly speaking, we should write $(\mu_{{\scriptscriptstyle\mathscr{X}}},X,\lambda)$ rationalizes $\rho$. However, since the role of $X$ is implicit in $\lambda$, we will simply write $(\mu_{{\scriptscriptstyle\mathscr{X}}},\lambda)$.
Note that the term $\mu_{{\scriptscriptstyle\mathscr{X}}}(\cdots)$ on the right-hand side represents the total market share of aggregate $a$ from menu $A$ within the subpopulation of consumers for whom each aggregate $b$ corresponds to the subset $S_{b}\subseteq X(b)$ of underlying alternatives. The term $\lambda_{A}((S_{b})_{b\in A})$ denotes the proportion of such consumers in the overall population. Thus, the right-hand side of equation (3) is a natural extension of the standard RUM: the observed choice probability is a weighted average of standard RUM probabilities, with weights given by $\lambda$ reflecting the distribution of compositions across consumers.
@@
Definition 2.4 assumes independence between the distribution $\mu_{{\scriptscriptstyle\mathscr{X}}}$ over preference rankings and the composition distribution $\lambda$ over underlying alternatives. While one could consider a more general framework allowing dependence between these components, such a model may lack testable implications, since the joint distribution could vary with the menu as $\lambda$ does. In particular, if the agent’s preferences may depend on composition, such dependence would allow the distribution over rankings to vary with the menu, rendering the model unfalsifiable. If, however, we only allow the composition distribution $\lambda$ to depend on the preferences, the characterization remains the same. See Remark D.2 in the online appendix for more details.
@@
ARU-rationality implies RU-rationality. To see why suppose $\rho$ is rationalized by $\mu_{{\scriptscriptstyle\mathscr{A}}}$ over aggregates $\mathscr{A}$. Choose an aggregation correspondence $X$ so that $X(a)$ is a singleton for all $a\in\mathscr{A}$. Label the sole element of $X(a)$ by $x_{a}$. Then for each $\succ$ in the support of $\mu_{{\scriptscriptstyle\mathscr{A}}}$ define $\succ^{\prime}$ by $x_{a}\succ^{\prime}x_{b}$ whenever $a\succ b$ and define all other comparisons arbitrary. By setting $\mu_{{\scriptscriptstyle\mathscr{X}}}(\succ^{\prime})=\mu_{{\scriptscriptstyle\mathscr{A}}}(\succ)$ for all $\succ$ in the support of $\mu_{{\scriptscriptstyle\mathscr{A}}}$, it is apparent that $(\mu_{{\scriptscriptstyle\mathscr{X}}},\lambda)$ rationalizes $\rho$.
@@
A stochastic choice function $\rho$ is partially RU-rational if there exists $\mu\in\Delta(\mathscr{L}(\mathscr{A}_{A}))$ such that for any $a\in A\subseteq\mathscr{A}_{A}$ such that $A\in\mathscr{D}$,
$\rho(A,a)=\mu(\succ\in\mathscr{L}(\mathscr{A}_{A})|a\succ b\text{ for all }b\in A\setminus a).$ (4)
@@
RU-rationality implies partial RU-rationality. When the menu contains only atomic aggregates, each aggregate maps to a single underlying alternative, so the RU-representation collapses to a standard RUM on the set of atomic aggregates, which corresponds to partial RU rationality; $\lambda$ terms disappear because there is no compositional uncertainty on such menus.
@@
As discussed in the introduction, RUM over the underlying alternatives is the natural assumption on the data generating process, while ARUM is the standard empirical tool. The central question is therefore whether choice frequencies generated by a RUM are also consistent with an ARUM.
@@ 3 Characterization
In this section, we characterize RU-rationality in three different frameworks and show precisely how its testable implications are weaker than those of ARU-rationality. Throughout this section except Section 3.3, we consider the outside option setup in which there exists only one non-atomic aggregate $a_{0}$. Formally, we set $\mathscr{A}_{N}=\{a_{0}\}$ and $\mathscr{A}_{A}=\mathscr{A}\setminus\{a_{0}\}$. We call $a_{0}$ the outside option. This case is useful for two reasons. First, it reflects the common modeling practice of including an outside option—an aggregate that captures unobserved or unlisted choices—frequently used in empirical applications. Second, this simplified setting serves as a foundation for analyzing the more general case in Theorem 3.5 with multiple non-atomic aggregates in Section 3.3.
We will make an assumption for the following theorems (Theorem 3.1 and Theorem 3.2) and its generalization (Theorem 3.5) for multiple non-atomic aggregates.
Assumption (Richness): The set $\mathscr{X}$ contains at least $|\mathscr{A}|^{2}$ distinct elements, so that each aggregate $a\in\mathscr{A}$ may be associated with an underlying set $X(a)$ containing at least $|\mathscr{A}|$ elements. [^10] This assumption is required for Theorems 3.1 and 3.5. Since the if direction of Theorem 3.2 relies on Theorem 3.1, it is also required there. For the outside option setup where $\mathscr{A}_{N}=\{a_{0}\}$, the bound can be weakened to $\mathscr{X}$ contains at least $2|\mathscr{A}_{A}|+1$ elements; see Theorem 3.3. The stronger $|\mathscr{A}|^{2}$ assumption is only needed for multiple non-atomic aggregates.
This richness assumption is natural in applications where one aggregate is the outside option: the outside option typically aggregates a large and heterogeneous collection of underlying alternatives. We study the case in which the aggregation correspondence is exogenously fixed and $\mathscr{X}$ is not sufficiently rich in Section 3.2.
@@
Assume Richness of $\mathscr{X}$. Consider the outside option setup (i.e., $\mathscr{A}_{N}=\{a_{0}\}$). A stochastic choice function $\rho$ is RU-rational if and only if $\rho$ satisfies the following two conditions:
@@
    (Limited Monotonicity): if $b\in D\subseteq\mathscr{A}_{A}$ then $\rho(D,b)\geq\rho(D\cup\{a_{0}\},b)$,
@@
The necessity of (i) and (ii) is immediate. The sufficiency part of the proof is constructive and nontrivial: given a stochastic choice function $\rho$ that satisfies the two conditions (i) and (ii), we need to construct an aggregation correspondence $X$, a composition distribution $\lambda$ and a distribution $\mu_{\mathscr{X}}$ over linear orders on $\mathscr{X}$ that satisfies equation (3). The proof is in the appendix.
The theorem makes precise the weakness of RU-rationality relative to ARU-rationality: RU-rationality is substantially weaker. To see this, note that Limited Monotonicity requires only that adding the outside option does not increase the choice frequencies of atomic aggregates. In particular, it does not require the monotonicity condition with respect to menus containing non-atomic aggregates. To see this, consider a choice set $D$ such that $a_{0}\in D$ and $b\notin D$. The theorem shows that under RU-rationality, it is possible to observe
@@
which violates the standard monotonicity condition. This may arise because, in the expanded set $D\cup{b}$, the outside option $a_{0}$ may become more desirable than it was in $D$. For example, in the context of cereal choice, the presence of a fancy, expensive cereal brand may signal that the market corresponds to a high-income area leading to a change in the interpretation or attractiveness of the outside good: the outside goods may contain more appealing options such as smoked salmon or quiche. Another example would be the newly added alternative $b$ is only available locally in a single area. Thus, $D\cup b$ corresponds to a special area which may contain a different composition of outside options. By contrast, the meanings of atomic aggregates remain the same when $a_{0}$ is added to a choice set; thus the limited monotonicity condition holds.
ARU-rationality, on the other hand, imposes much stronger restrictions including full monotonicity and beyond. When $\rho$ is defined on all subsets of $\mathscr{A}$ (i.e., $\mathscr{D}=2^{\mathscr{A}}\setminus\{\emptyset\}$), ARU-rationality is equivalent to the non-negativity of all Block-Marschak (BM) polynomials. Among these, monotonicity is one of the simplest and most intuitive implications. However, as Theorem 3.1 shows, RU-rationality only implies monotonicity on limited choice sets. This result highlights the substantial gap between RU-rationality and ARU-rationality.
A few more technical remarks are in order:
@@
In our discussion, we focus on Limited Monotonicity rather than on partial RU-rationality. The reason is straightforward: RU-rationality on atomic alternatives is well studied. [^11] For example, under the full domain assumption (i.e., $\mathscr{D}=2^{\mathscr{A}}\setminus\{\emptyset\}$), it can be characterized by the non-negativity of Block–Marschak (BM) polynomials defined over sets composed solely of atomic aggregates.
@@
In our notion of RU rationality, we assume that the composition of aggregates and the agent’s preferences are independent. A natural extension is to allow the composition distribution to depend on the agent’s preferences. In Remark D.2 of the online appendix, we show such a representation (3) is also characterized by Limited Monotonicity and Partial-RU rationality.
@@
Readers may wonder about the relationship between RU-rationality and the nested logit model. First, nested logit does not address the fundamental problem that the economic meaning of the outside option may vary across choice sets. Second, even when the composition $X(a_{0})$ is known, nested logit is most effective precisely in environments where the underlying alternatives within the outside option share highly correlated utility shocks—as in the canonical red-bus/blue-bus example (see Section 4.2.2 of Train, 2003). In Section 4.1, we show that this case corresponds exactly to a condition called non-overlappingness of preferences under which RU-rationality implies ARU-rationality. In such settings, ARUM may entail little additional bias, so nested logit offers no real advantage.